Prove the 3√7 is irrational 


1
Expert's answer
2022-05-15T12:16:26-0400

Let "3\sqrt7" be a rational Number in form of "\frac p q" where "p" , "q" are coprimes:

"3\sqrt7=\frac p q"

Square on both sides:

"9*7=\frac {p^2} {q^2}"

"63q^2=p^2"

Since "p" and "q" are coprimes we can say that 63 is a factor of "p^2", then "p" is also a factor of 63 because it is a rational number.

Hence "p" can be expressed as "k" times 63 where k is some constant:

"p=63k"

Then "63q^2=(63k)^2"

"\implies q^2=63k^2"

From here we get

63 is a factor of "q^2" and 63 is a factor of "q".

We got that 63 is a factor of "p" and "q".

As we assumed "p" and "q" are coprimes and it proves our assumption was wrong and nereby we can say "3\sqrt7" is an irrational number.